# reference.bib

### v2

@@ II MONOCULDAR 3D LANE DETECTION IN AD
Figure 3: The Gen-LaneNet procedure involves encoding an input image into deep features using a segmentation backbone, which are then decoded into a lane segmentation map. The 3D-GeoNet then focuses on geometry encoding and predicts intermediate 3D lane points represented in top-view 2D coordinates and real heights. Finally, a geometric transformation converts the network output into real-world 3D lane points.
Liu et al. proposed CLGo [liu2022learning], a two-stage framework for predicting 3D lanes and camera pose from a single image. Its first stage focuses on camera pose estimation and introduces an auxiliary 3D lane task and geometry constraints for multi-task learning. The second stage targets the 3D lane task and uses the previously estimated pose to generate top-view images for accurate 3D lane prediction. PersFormer [chen2022persformer] introduces the first transformer-based approach for 3D lane detection and proposes a novel architecture called Perspective Transformer, as shown in Fig. 4. This transformer-based architecture enables spatial feature transformation, allowing for accurate detection of 3D lane lines. Moreover, the proposed framework has the unique capability to simultaneously handle both 2D and 3D lane detection tasks, providing a unified solution. Additionally, the paper presents OpenLane, a large-scale 3D lane detection dataset built upon the influential Waymo Open dataset [sun2020scalability]. OpenLane is the first dataset of its kind to offer high-quality annotations and diverse real-world scenarios, making it a valuable resource for advancing research in the field.
Figure 4: PersFormer pipeline. The key is to understand the spatial feature transformation from the front view to the BEV space, aiming to enhance the representativeness of the resulting BEV features at the target point by taking into account the local context around the reference point.
In [yan2022once], researchers introducesd the largest real-world 3D lane detection dataset, ONCE-3DLanes dataset, and provides a more comprehensive evaluation metric to rekindle interest in this task in real-world scenarios. Additionally, this paper proposes a method called SALAD, which enables direct generation of 3D lane layouts from front view images without the need for converting the feature map into the bird’s-eye view (BEV), the network architecture of SALAD is shown in Fig. 5. This paper [li2022reconstruct] proposes a novel loss function that utilizes the geometry structure prior of lanes in 3D space, enabling stable reconstruction from local to global and providing explicit supervision. It introduces a 2D lane feature extraction module that leverages direct supervision from the top view, ensuring maximum retention of lane structural information, especially in distant areas, the overrall pipeline is shown is Fig. 7. Additionally, the paper presents a task-specific data augmentation method for 3D lane detection to address the imbalanced data distribution of ground slope and camera pose, enhancing generalization on rare cases. Bai et al. proposed CurveFormer [bai2022curveformer], a Transformer-based algorithm for 3D lane detection. In this paper, researchers formulate the queries in the decoder layers as a dynamic anchor point set, and utilize a curve cross-attention module to compute the similarity between the queries and image features. Besides, they also introduce a context sampling unit to predict offsets by combining reference features and queries, guiding the learning of sampling offsets. Ai et al. proposed WS-3D-Lane [ai2022ws], which for the first time proposed a weakly supervised 3D lane detection method using only 2D lane labels, and outperformed the previous 3D-LaneNet [garnett20193d] method in evaluation. In addition, the author proposed a camera pitch self-calibration method, which can calculate the camera pitch angle online in real time, thereby reducing the error caused by the pitch angle change between the camera and the ground plane caused by the uneven road surface. In BEV-LaneDet [wang2022bev], the author introduced the concept of a virtual camera, a novel module for data prediction processing, aiming to standardize the extrinsic parameters of cameras and ensure the consistency of data distribution. Furthermore, they proposed the use of keypoint representation as a simple yet effective method for representing 3D lane structures. Additionally, they presented the MLP Based Spatial Transformation Pyramid, a lightweight architecture designed to convert multi-angle visual features into BEV features. Huang et al. proposed the Anchor3DLane [huang2023anchor3dlane] framework, which directly defines anchor points in 3D space and directly regresses 3D lanes from the front view without introducing a top view, as shown in Fig. 6. The authors also proposed a multi-frame extension of Anchor3DLane to exploit well-aligned temporal information and further improve performance. Furthermore, a global optimization method is developed to fine-tune the lane lines by utilizing the lane iso-width property.
Figure 5: The architecture of SALAD. The input image is encoded into deep features by the backbone, and then two branches, the semantic awareness branch and the spatial contextual branch, decode the features to extract the spatial information of the lane and create the segmentation mask. Subsequently, the 3D reconstruction process integrates this information to ultimately derive the 3D lane positions in the real-world scene.
@@
Figure 6: The architecture of Anchor3DLane. Utilizing a front-view input image, we employ a CNN backbone and a Transformer layer to extract the visual feature F initially. Subsequently, 3D anchors are projected to sample their features from F based on camera parameters. Following this, a classification head and a regression head are utilized to make the final predictions. The lane predictions can also be used as new 3D anchors for iterative regression.
@@
Figure 7: The framework includes a top-view segmentation network and an anchor-based 3D lane regressor. Features from a front-view image are projected onto the top view to create the segmentation mask under top-view supervision. The anchor representation of 3D lanes is estimated from the top-view lane mask. Additionally, we incorporate the geometry prior from 3D to 2D projection for 3D lane reconstruction.
@@
Li et al. proposed a method [li2022reconstruct] to directly extract top view lane information from front-view images, reducing the structural loss in the 2D lane representation. The overall pipeline of this method is shown in the Fig. 7. In this paper, the authors considered 3D lane detection as a reconstruction problem from 2D images to 3D space. They proposed that explicit imposition of the geometric prior of 3D lanes during the training process is essential for fully utilizing the structural constraints of inter-lane and intra-lane relationships, and for extracting the height information of 3D lanes from the 2D lane representation. The authors analyzed the geometric relationship between 3D lanes and their 2D representation, and proposed an auxiliary loss function based on the geometric structure prior. They also demonstrated that explicit geometric supervision can enhance noise elimination, outlier rejection, and structure preservation for 3D lanes. Bai et al. proposed CurveFormer [bai2022curveformer] and CurveFormer++ [bai2024curveformer++], which are single-stage methods based on Transformer that can directly calculate the parameters of 3D lanes and can bypass the challenging view transformation step. Specifically, they formalized 3D lane detection as a curve propagation problem using curve queries. A 3D lane query is represented by a dynamic and ordered set of anchor points. By using queries with curve representation in the Transformer decoder, the 3D lane detection results are iteratively refined. Additionally, they introduced a curve cross-attention module to calculate the similarities between curve queries and image features. Furthermore, a context sampling module is provided to capture more relevant image features of a curve query, further enhancing the performance of 3D lane detection. Similar to [qin2020ultra], Li et al. proposed GroupLane [li2023grouplane], a row-wise classification-based method for 3D lane detection. GroupLane’s design consists of two groups of convolutional heads, each corresponding to a lane prediction. This grouping decouples the interaction of information among different lanes, reducing optimization difficulty. During training, predictions are matched with lane labels using a single-win one-to-one matching (SOM) strategy, which assigns predictions to the most suitable labels for loss computation. To address the challenge posed by the inevitable depth ambiguity in monocular images causing misalignment between the constructed surrogate feature map and the original image for lane detection, Luo et al. proposed a novel LATR model [luo2023latr]. This is an end-to-end 3D lane detector that uses 3D-aware front-view features without the need for transformed view representation. Specifically, LATR detects 3D lanes through cross-attention based on query and key-value pairs constructed using a lane-aware query generator and dynamic 3D ground positional embedding. On one hand, each query is generated based on 2D lane-aware features and adopts a hybrid embedding to enhance lane information. On the other hand, 3D spatial information is injected as positional embedding from an iteratively updated 3D ground plane. To address the issue of inaccurate view transformation caused by the neglect of road height changes in the process of transforming image-view features into bird’s-eye view using inverse perspective mapping (IPM), Chen et al. proposed an efficient transformer for 3D lane detection [chen2023efficient]. Unlike conventional transformers, the model includes a decomposed cross-attention mechanism that can simultaneously learn lane and bird’s-eye view representations. This approach allows for more accurate view transformation compared to IPM-based methods and is more efficient. Previous studies assumed that all lanes are on a flat ground plane. However, Kim et al. argued that algorithms based on this assumption have difficulty in detecting various lanes in actual driving environments and proposed a new algorithm, D-3DLD [kim2023d]. Unlike previous methods, this approach expands rich contextual features from the image domain to 3D space by utilizing depth-aware voxel mapping. Additionally, the method determines 3D lanes based on voxelized features. The authors designed a new lane representation combined with uncertainties and predicted the confidence intervals of 3D lane points using Laplace loss. Li et al. proposed a lightweight method [li2022perspective] that uses MobileNet [howard2017mobilenets] as the backbone network to reduce the demand for computational resources. The proposed method includes the following three stages. First, the MobileNet model is used to generate multi-scale front-view features from a single RGB image. Then, a perspective transformer calculates the bird’s eye view (BEV) features from the front-view features. Finally, two convolutional neural networks are used to predict the 2D and 3D coordinates and their respective lane types. In paper [han2023decoupling], Han et al. argued that the curve-based lane representation may not be well-suited for many irregular lane lines in real-world scenarios, which can lead to performance gaps compared to indirect representations such as segmentation-based or point-based methods. In this paper, the authors proposed a new lane detection method, which can be decomposed into two parts: curve modeling and ground height regression. Specifically, a parameterized curve is used to represent lanes in the bird’s eye view space to reflect the original distribution of lanes. For the second part, since ground heights are determined by natural factors such as road conditions and are not comprehensive, the ground heights are regressed separately from the curve modeling. Additionally, the authors have designed a new framework and a series of losses to unify the 2D and 3D lane detection tasks, guiding the optimization of models with or without 3D lane labels.
@@
An intuitive summary of these methods is shown in Table LABEL:tab:overviw, which includes a description of the methods, the datasets used, the open-source status, and the network architecture.
Figure 8: Several open-source algorithms’ visualization results on the ApolloSim dataset, with 3D-LaneNet, Gen-LaneNet, CLGo, WS-3D-Lane, and Anchor-3D-Lane algorithms displayed from top to bottom. In the images, the red lines represent the ground truth of the lane lines, while the blue lines represent the lane lines predicted by the algorithms.
TABLE I: LIST OF USED VARIABLES
Variable Description
@@
$L^{p}$ The predicted lane.
$L^{g}$ The ground-truth lane.
$P_{gi}$ Equals to $(x_{gi},y_{gi},z_{gi})$, a point of $L^{g}$.
$P_{pj}$ Equals to $(x_{pj},y_{pj},z_{pj})$, a point of $L^{p}$.
@@
$\gamma$ TA threshold parameter in Huber loss, defining the point where the loss function transitions from quadratic to linear.

### v3

@@ I-D Review Methodology
Figure 3: The procedures of Gen-LaneNet [guo2020gen] involve encoding an input image into deep features using a segmentation backbone, which are then decoded into a lane segmentation map. The 3D-GeoNet then focuses on geometry encoding and predicts intermediate 3D lane points represented in top-view 2D coordinates and real heights. Finally, a geometric transformation converts the network output into real-world 3D lane points.
II MONOCULAR 3D LANE DETECTION IN AD
With the rapid development of AD and deep learning, learning-based monocular lane detection has received increasing attention from both the industry and academia. In the field of monocular lane detection, early work mainly focused on 2D lane detection. As AD technology matures and the demand for cost reduction increases, higher requirements are placed on lane detection, aiming to predict the 3D information of lane lines from a single image.Therefore, starting from 2018, there have been successive works on monocular 3D lane detection. As shown in Fig. 1, the figure provides a chronological overview of the monocular 3D lane detection algorithms. It can be seen that with the passage of time, more and more works have emerged, indicating that this field is becoming increasingly popular. In this figure, the green arrows represent CNN-based methods, the orange arrows represent Transformer-based methods, and the blue arrows represent methods with hybrid architecture.
Among these methods, 3D-LaneNet [garnett20193d] is a pioneering work in the field of monocular 3D lane detection. The overview of 3D-LaneNet is depicted in the Fig. 2. 3D-LaneNet introduces a network that can directly predict 3D lane information in road scenes from monocular images. This work is the first to solve the task of 3D lane detection using on-board monocular vision sensor. 3D-LaneNet introduces two novel concepts: intra-network feature map IPM and anchor-based lane representation. The intra-network IPM projection facilitates dual representation information flow in both front view and bird’s-eye view. The anchor-based lane output representation supports an end-to-end training approach, which has demonstrated a significant impact on subsequent work in this field. Many later studies have adopted the anchor-based lane representation. Although 3D-LaneNet has shown promising results in 3D lane detection, its limitations lie in using an inappropriate coordinate frame that misaligns lane points with visual features, and in the end-to-end learning framework where geometric encoding is easily affected by changes in image appearance, thus requiring more training data, which limits its practical application. Inspired by 3D-LaneNet and its drawbacks, Guo et al. propose Gen-LaneNet [guo2020gen], a generalized and scalable framework for 3D lane detection, as shown in Fig. 3. Compared to 3D-LaneNet, Gen-LaneNet is still a unified framework that addresses image encoding, spatial transformation of features, and 3D curve extraction within a single network. However, it has major differences in two aspects: a geometric extension to the lane anchor design and a scalable two-stage network that decouples the learning of image encoding and 3D geometry reasoning. Additionally, this paper introduces a highly realistic synthetic image dataset with rich visual variations for developing and evaluating 3D lane detection methods. Although the new geometry-guided lane anchor representation has better generalization ability to unobserved scenes, it is still limited to long lanes that are roughly parallel to the direction of travel of the ego vehicle. Based on Gen-LaneNet, Jiang et al. propose Att-Gen-LaneNet [jiang2022robust], which incorporates two attention mechanisms: the Efficient Channel Attention (ECA) [wang2020eca] attention mechanism and the Convolutional Block Attention Module (CBAM) [woo2018cbam] attention mechanism to enhance algorithm performance, ultimately achieving an improvement in performance compared to Gen-LaneNet. Subsequently, Efrat et al. propose a tile representation to expand support for lane topologies. In the two works proposed by Efrat et al. [efrat2020semi, efrat20203d], a semi-local tile-based representation is employed, which captures local lane structures and road geometrical features, along with a learning-based method that clusters local lane segments into complete 3D lane curves. The difference is that in [efrat2020semi], uncertainty estimation is introduced into 3D lane line detection for the first time to reflect the detection noise. Such methods still have certain limitations when dealing with complex road curvatures and geometries. Despite using a semi-local tile representation, performance may degrade in handling some extreme cases, such as highly curved roads. Instead of utilizing the ground truth camera pose, Liu et al. propose CLGO [liu2022learning], a hybrid architecture method that combines CNN and Transformer, can learn camera pose online. This method follows a two-stage framework. In the first stage, camera pose is estimated by introducing an auxiliary 3D lane task and geometric constraints to improve pose estimation. In the second stage, the estimated camera pose is used to generate top-view images, which are then used for accurate 3D lane prediction. It is worth mentioning that, unlike anchor-based representations, this method represents lane lines using polynomials. The advantage of this method is that it can estimate camera parameters in real-time and reduce errors. However, the drawback is that it introduces an additional camera pose estimation step, which increases the system’s complexity. Another hybrid architecture method is the Anchor3DLane [huang2023anchor3dlane] framework, which directly defines anchor points in 3D space and directly regresses 3D lanes from the front view without introducing a top view, as shown in Fig. 4. The authors also propose a multi-frame extension of Anchor3DLane to exploit well-aligned temporal information and further improve performance. Furthermore, a global optimization method is developed to fine-tune the lane lines by utilizing the lane iso-width property.
@@ II MONOCULAR 3D LANE DETECTION IN AD
Figure 4: The architecture of Anchor3DLane [huang2023anchor3dlane]. Utilizing a front-view input image, we employ a CNN backbone and a Transformer layer to extract the visual feature $F$ initially. Subsequently, 3D anchors are projected to sample their features from $F$ based on camera parameters. Following this, a classification head and a regression head are utilized to make the final predictions. The lane predictions can also be used as new 3D anchors for iterative regression.
Figure 5: The pipeline of PersFormer [chen2022persformer]. The key is to understand the spatial feature transformation from the front view to the BEV space, aiming to enhance the representativeness of the resulting BEV features at the target point by taking into account the local context around the reference point.
Figure 6: The architecture of SALAD [yan2022once]. The input image is encoded into deep features by the backbone, and then two branches, the semantic awareness branch and the spatial contextual branch, decode the features to extract the spatial information of the lane and create the segmentation mask. Subsequently, the 3D reconstruction process integrates this information to ultimately derive the 3D lane positions in the real-world scene.
The above methods are all CNN-based approaches. PersFormer [chen2022persformer] introduces the first Transformer-based approach for 3D lane detection and proposes a novel architecture called Perspective Transformer, as shown in Fig. 5. This Transformer-based architecture enables spatial feature transformation, allowing for accurate detection of 3D lane lines. Moreover, the proposed framework has the unique capability to simultaneously handle both 2D and 3D lane detection tasks, providing a unified solution. This method not only achieved the best performance at the time in the 3D lane detection task but also performes well in 2D lane detection. However, compared to previous methods, its inference time is longer, which slightly affects real-time performance. Additionally, the work presents OpenLane, a large-scale 3D lane detection dataset built upon the influential Waymo Open dataset [sun2020scalability]. OpenLane is the first dataset of its kind to offer high-quality annotations and diverse real-world scenarios, making it a valuable resource for advancing research in the field. To avoid converting the feature map into a BEV image, Yan et al. propose an extrinsic-free, anchor-free method called SALAD [yan2022once], as shown in Fig. 6, which can directly regress the 3D information of lanes in the image view. Specifically, the lane lines are segmented in the front view, and then the depth information is used to recover the 3D information of the lanes. It suffers from certain limitations, since this method relies on LiDAR to provide 3D information for training, issues such as occlusion or point cloud sparsity at a distance can affect the algorithm’s performance. Additionally, this paper introducesd the largest real-world 3D lane detection dataset, ONCE-3DLanes dataset, and provides a more comprehensive evaluation metric to rekindle interest in this task in real-world scenarios.
@@
Figure 7: The architecture of [li2022reconstruct]. The framework includes a top-view segmentation network and an anchor-based 3D lane regressor. Features from a front-view image are projected onto the top view to create the segmentation mask under top-view supervision. The anchor representation of 3D lanes is estimated from the top-view lane mask. Additionally, the geometry prior from 3D to 2D projection is incorporated for 3D lane reconstruction.
Li et al. [li2022reconstruct] believe that the structure of 3D lanes and the 2D-3D relationships following geometric priors should be jointly optimized throughout the entire process of 3D lane detection. They propose a novel loss function that utilizes the geometry structure prior of lanes in 3D space, enabling stable reconstruction from local to global and providing explicit supervision. It introduces a 2D lane feature extraction module that leverages direct supervision from the top view, ensuring maximum retention of lane structural information, especially in distant areas, the overall pipeline is shown in Fig. 7. There are also two Transformer-based developments, CurveFormer [bai2022curveformer] and CurveFormer++ [bai2024curveformer++], with CurveFormer++ being an improved version of CurveFormer. Their core idea is to represent lanes as a series of ordered 3D points, known as curve queries, and to iteratively optimize these queries in the Transformer decoder to accurately detect 3D lane lines. The improvements in CurveFormer++ involve incorporating temporal information fusion and a dynamic anchor set, as well as optimizing the network structure and cross-attention mechanism, achieving higher performance and better generalization capabilities in the 3D lane detection task. However, despite the progress made by CurveFormer and CurveFormer++ in the field of 3D lane detection, they also have certain drawbacks, such as high computational resource requirements and longer inference times.
@@
Due to the challenge posed by different cameras with varying parameters in 3D lane detection, Wang et al. introduce the concept of a Virtual Camera in BEV-LaneDet [wang2023bev] to unify these parameters. In addition, the method proposes a simple yet effective 3D lane representation called Key-Points Representation, which is better suited for representing complex and diverse 3D lane structures. Moreover, a lightweight and chip-friendly Spatial Transformation Pyramid module was designed to convert multiscale front-view features into BEV features. This approach achieves excellent performance while maintaining fast inference speed. However, a drawback is that since this method focuses more on the BEV plane, its performance in terms of $z$-axis error is not ideal, requiring further improvement. Li et al. propose a method [li2022reconstruct] to directly extract top view lane information from front-view images, reducing the structural loss in the 2D lane representation. The overall pipeline of this method is shown in the Fig. 7. In this work, the authors consider 3D lane detection as a reconstruction problem from 2D images to 3D space. They propose that explicit imposition of the geometric prior of 3D lanes during the training process is essential for fully utilizing the structural constraints of inter-lane and intra-lane relationships, and for extracting the height information of 3D lanes from the 2D lane representation. The authors analyze the geometric relationship between 3D lanes and their 2D representation, and propose an auxiliary loss function based on the geometric structure prior. They also demonstrate that explicit geometric supervision can enhance noise elimination, outlier rejection, and structure preservation for 3D lanes. Similar to [qin2020ultra], Li et al. propose GroupLane [li2023grouplane], a row-wise classification-based method for 3D lane detection. GroupLane’s design consists of two groups of convolutional heads, with each corresponding to a lane prediction. This grouping decouples the interaction of information among different lanes, reducing the difficulty in optimization. During training, predictions are matched with lane labels using a single-win one-to-one matching strategy, which assigns predictions to the most suitable labels for loss computation. To address the challenge posed by the inevitable depth ambiguity in monocular images causing misalignment between the constructed surrogate feature map and the original image for lane detection, Luo et al. propose a novel LATR model [luo2023latr]. This is an end-to-end 3D lane detector that uses 3D-aware front-view features without the need for transformed view representation. Specifically, LATR detects 3D lanes through cross-attention based on query and key-value pairs constructed using a lane-aware query generator and dynamic 3D ground positional embedding. On one hand, each query is generated based on 2D lane-aware features and adopts a hybrid embedding to enhance lane information. On the other hand, 3D spatial information is injected as positional embedding from an iteratively updated 3D ground plane. To address the issue of inaccurate view transformation caused by the neglect of road height changes in the process of transforming image-view features into BEV using IPM, Chen et al. propose an efficient transformer for 3D lane detection [chen2023efficient]. Unlike conventional transformers, the model includes a decomposed cross-attention mechanism that can simultaneously learn lane and BEV representations. This approach allows for more accurate view transformation compared to IPM-based methods and is more efficient. Previous studies assume that all lanes are on a flat ground plane. However, Kim et al. argue that algorithms based on this assumption have difficulty in detecting various lanes in actual driving environments and propose a new algorithm, D-3DLD [kim2023d]. Unlike previous methods, this approach expands rich contextual features from the image domain to 3D space by utilizing depth-aware voxel mapping. Additionally, the method determines 3D lanes based on voxelized features. The authors design a new lane representation combined with uncertainties and predicted the confidence intervals of 3D lane points using Laplace loss. Li et al. propose a lightweight method [li2022perspective] that uses MobileNet [howard2017mobilenets] as the backbone network to reduce the demand for computational resources. The proposed method includes the following three stages. First, the MobileNet model is used to generate multi-scale front-view features from a single RGB image. Then, a perspective Transformer calculates the BEV features from the front-view features. Finally, two convolutional neural networks are used to predict the 2D and 3D coordinates and their respective lane types. In paper [han2023decoupling], Han et al. argue that the curve-based lane representation may not be well-suited for many irregular lane lines in real-world scenarios, which can lead to performance gaps compared to indirect representations such as segmentation-based or point-based methods. In this paper, the authors propose a new lane detection method, which can be decomposed into two parts: curve modeling and ground height regression. Specifically, a parameterized curve is used to represent lanes in the BEV space to reflect the original distribution of lanes. For the second part, since ground heights are determined by natural factors such as road conditions and they are not comprehensive, the ground heights are regressed separately from the curve modeling. Additionally, the authors have designed a new framework and a series of losses to unify the 2D and 3D lane detection tasks, guiding the optimization of models with or without 3D lane labels.
The above methods only consider single-frame input, which leads to poor performance in scenarios without visual cues (e.g., obscured by surrounding vehicles). To address this issue and better exploit the spatio-temporal continuity of lanes, Wang et al. propose STLane3D [wang2022spatio]. This approach leverages a multi-frame fusion mechanism that utilizes the spatio-temporal continuity of lanes to resolve the problem of insufficient visual cues in single-frame detection. Additionally, they introduce a pre-alignment operation and a spatio-temporal attention module to effectively fuse multi-frame features, as well as a 3DLane IOULoss, which constrains the lateral and longitudinal changes of the predicted lanes to improve performance. The method achieves state-of-the-art results on public datasets.
Learning-based methods cannot avoid relying on training data, however, annotating this training data is a high-cost and labor-intensive task, which hinders the application of 3D lane detection. In light of this, some researchers have proposed weakly supervised and self-supervised methods to reduce dependence on manually labeled data. Ai et al. propose a weakly supervised method called WS-3D-Lane [ai2022ws], which for the first time proposes a weakly supervised 3D lane detection method using only 2D lane labels, and outperform the previous 3D-LaneNet [garnett20193d] method in evaluation. In addition, the authors propose a camera pitch self-calibration method, which can calculate the camera pitch angle online in real time, thereby reducing the error caused by the pitch angle change between the camera and the ground plane caused by the uneven road surface. Although this method reduces the annotation requirements for 3D lane lines, it still relies on manually labeled 2D lane line data, and the assumption of constant lane width does not hold in some scenarios, which may affect the performance of the method. To further reduce the reliance on manual data annotation, Nie et al. propose LaneCorrect [nie2024lanecorrect], a self-supervised lane detection method. This approach introduces an unsupervised 3D lane segmentation technique that leverages the unique intensity of lanes in LiDAR point clouds to generate noisy 2D lane labels. Additionally, it presents a self-supervised training scheme that automatically corrects noisy lane labels by learning geometric consistency and instance-aware adversarial augmentation. Finally, the method introduces a distillation technique that allows the student network to be trained for any target lane detection dataset without using any labeled data. Although this method significantly reduces the reliance on manual data annotation, the generated pseudo-labels contain some noises, such as projection errors and mislabeling caused by clustering errors, which can affect its performance.
@@
An intuitive summary of these methods is shown in Table I, which includes a description of the methods, the datasets used, the open-source status, and the network architecture.
TABLE I: A detailed description of existing monocular 3D lane detection methods.
Ref. Year Description Dataset Code Lane Representation Network Architecture   Need BEV Trans.?
[garnett20193d] 2018 This work proposes a method that combines a novel network architecture, IPM projections, and anchor-based lane representation to achieve accurate 3D lane detection. Synthetic-3D-Lanes, 3D-Lanes N/A Anchor representation           CNN                 ✔
[guo2020gen] 2020 This work employs a two-stage framework that decouples the learning of the image segmentation subnetwork from the geometry encoding subnetwork, thereby reducing the amount of 3D lane labels required for training. Apollo Synthetic. link Anchor representation           CNN                 ✔
[efrat2020semi] 2020 This approach introduces uncertainty estimation, utilizes a semi-local BEV tile representation to decompose lanes into simple segments, and subsequently clusters these segments into complete lanes. Synthetic-3D-Lanes, 3D-lanes. N/A Tile representation           CNN                 ✔
[efrat20203d] 2020 This method is anchor-free and capable of detecting complex lane topologies, including splits, merges, short lanes, and perpendicular lanes. Synthetic-3D-Lanes, Real-3D-lanes N/A Tile representation           CNN                 ✔
[liu2022learning] 2021 This work presents a two-stage framework that first leverages auxiliary 3D lane tasks and geometry constraints, and then performs accurate 3D lane prediction in the second stage using the estimated pose. Apollo Synthetic. link Polynomial representation         Hybrid                ✔
[chen2022persformer] 2022 This paper introduces a novel Transformer-based spatial feature transformation module and presents OpenLane, a large-scale real-world 3D lane dataset. Apollo Synthetic, OpenLane. link Anchor representation        Transformer           ✔
[yan2022once] 2022 This paper introduces the ONCE-3DLanes dataset and proposes SALAD, an extrinsic-free, anchor-free method for regressing 3D lane coordinates in image view. ONCE-3DLanes. link Pixel representation           CNN                 ✘
[jiang2022robust] 2022 This paper proposes Att-Gen-LaneNet, a two-stage network for robust 3D lane detection. It integrates the Efficient Channel Attention (ECA) and Convolutional Block Attention Module (CBAM) to improve segmentation and generalization. Apollo Synthetic. N/A Anchor representation           CNN                 ✔
[li2022reconstruct] 2022 This paper proposes an approach for monocular 3D lane detection by leveraging the geometry structure and explicit supervision based on the structure prior. Apollo Synthetic. N/A Anchor representation           CNN                 ✔
[li2022perspective] 2022 This method comprises three stages: first, generating multiscale front-view features; second, calculating BEV features; and third, predicting both 2D and 3D lanes. OpenLane. N/A Anchor representation           CNN                 ✔
[bai2022curveformer] 2022 This method formulates 3D lane detection as a curve propagation problem, enhancing performance by leveraging curve queries, a curve cross-attention module, and a context sampling module to capture relevant image features. Apollo Synthetic, OpenLane. N/A Anchor representation        Transformer           ✘
[ai2022ws] 2022 This paper introduces a method training 3D lanes with only 2D lane labels, leveraging assumptions of constant lane width and equal height on adjacent lanes for indirect supervision of 3D lane heights. Apollo Synthetic, ONCE-3DLanes. N/A Anchor representation           CNN                 ✔
[wang2023bev] 2022 This paper introduces a Virtual Camera for spatial relationship consistency, a Key-Points Representation module for diverse 3D lane structures, and a Spatial Transformation Pyramid for front-view to BEV feature transformation. Apollo Synthetic, OpenLane. link Key-points representation           CNN                 ✔
[wang2022spatio] 2022 This model utilizes multi-frame input and removes dependency on camera pose by integrating a multi-frame pre-alignment layer, a spatio-temporal attention module, and 3DLane IOULoss. OpenLane, ONCE-3DLanes. N/A Anchor representation           CNN                 ✔
[huang2023anchor3dlane] 2023 This paper introduces Anchor3DLane, a BEV-free method predicting 3D lanes directly from front-viewed (FV) representations using defined 3D lane anchors. Apollo Synthetic, OpenLane, ONCE-3DLanes. link Anchor representation         Hybrid                ✘
[kim2023d] 2023 This paper introduces a method that employs depth-aware voxel mapping, extends contextual features to 3D space, and predicts confidence intervals using Laplace loss. LLAMAS-3DLD, Pandaset-3DLD, Apollo Synthetic. N/A Anchor representation           CNN                 ✔
[chen2023efficient] 2023 This paper proposes an efficient transformer for 3D lane detection using a decomposed cross-attention mechanism to learn lane and BEV representations concurrently. OpenLane, ONCE-3DLanes. N/A Key-points representation        Transformer           ✘
[li2023grouplane] 2023 This work, similar to [qin2020ultra], pioneers 3D lane detection by utilizing fully convolutional heads based on row-wise classification. OpenLane, ONCE-3DLanes, OpenLane-Huawei. N/A Row-wise representation           CNN                 ✔
[luo2023latr] 2023 This paper introduces LATR, an end-to-end 3D lane detection model utilizing 3D-aware front-view features and cross-attention mechanism with lane-aware query generator. OpenLane, Apollo Synthetic, ONCE-3DLanes. link Key-points representation Transformer            ✘
[han2023decoupling] 2023 This paper introduces a novel lane detection method, dividing it into curve modeling and ground height regression, within a unified framework for both 2D and 3D lane detection. OpenLane, ONCE-3DLanes, TuSimple, CULane. N/A Anchor representation           CNN                 ✘
[bai2024curveformer++] 2024 This approach models the 3D detection task as a curve propagation problem, where each lane is represented by a curve query with a dynamic and ordered anchor point set. OpenLane, ONCE-3DLanes. N/A Anchor representation Transformer            ✘
[nie2024lanecorrect] 2024 This paper introduces a self-supervised lane detection method that uses LiDAR intensity to generate noisy lane labels. Through geometric consistency and adversarially enhanced instance awareness, it corrects these labels without requiring manually annotated data. OpenLane, ONCE-3DLanes. N/A Anchor representation           CNN                 ✘
TABLE II: LIST OF USED VARIABLES
Notation Description
@@
$P$ A set of prediction lane points.
$Q$ A set of ground truth lane points.
$p$ A point of $P$.
$q$ A point of $Q$.
@@
$\gamma$ Threshold parameter in Huber loss, defining the point where the loss function transitions from quadratic to linear.